HNO 3 used as a reagent in stoichiometric analysis has specific gravity of 1.5 and contains 63% HNO 3 by mass. The volume of magnesium hydroxide solution which is 2.9 g/L in Mg(OH) 2 required for neutralizing 20 mL of above HNO 3 solution is?


HNO3 used as a reagent in stoichiometric analysis has specific gravity of 1.5 and contains 63% HNO3 by mass. The volume of magnesium hydroxide solution which is 2.9 g/L in Mg(OH)2 required for neutralizing 20 mL of above HNO3 solution is?


Grade:11

1 Answers

Souradip Sanyal
12 years ago

2HNO3+Mg(OH)2=Mg(NO3)2+2H2O.

Mass of HNO3 taken is 1.5*20/1000 gm = 0.03 gm.

Actual mass of HNO3  present 63% of 0.03 gm = 0.0189 gm = 0.0189/63 moles = 0.0003 mole.

Moles of H+ present = 0.0003 .

Moles of OH- needed (from Mg(OH)2) = 0.0003.

So, Moles of Mg(OH)2 needed = 0.00015 = 0.008745 gm Mg(OH)2 .

Let volume needed be v mL.

Then, 2.9*v/1000 = 0.008745

=> v = 3.01.

Ans:3.01 g of 100% pure Mg(OH)2 solution will be needed.